# RadFlag: A Black-Box Hallucination Detection Method for Medical Vision Language Models

### v1

@@ Abstract
Generating accurate radiology reports from medical images is a clinically important but challenging task. While current Vision Language Models (VLMs) show promise, they are prone to generating hallucinations, potentially compromising patient care. We introduce RadFlag, a black-box method to enhance the accuracy of radiology report generation. Our method uses a sampling-based flagging technique to find hallucinatory generations that should be removed. We first sample multiple reports at varying temperatures and then use a Large Language Model (LLM) to identify claims that are not consistently supported across samples, indicating that the model has low confidence in those claims. Using a calibrated threshold, we flag a fraction of these claims as likely hallucinations, which should undergo extra review or be automatically rejected. Our method achieves high precision when identifying both individual hallucinatory sentences and reports that contain hallucinations. As an easy-to-use, black-box system that only requires access to a model’s temperature parameter, RadFlag is compatible with a wide range of radiology report generation models and has the potential to broadly improve the quality of automated radiology reporting.
@@
The primary dataset used in this research is MIMIC-CXR, an open-source chest X-ray report generation dataset, publicly available through PhysioNet (Johnson et al., 2019). We plan to make our code publicly available.
@@ 1 Introduction
Generating accurate radiology reports from medical images is a challenging task that requires a deep understanding of medical imaging and the ability to precisely interpret and communicate complex findings. Medical Vision Language Models (VLMs) that automatically generate radiology reports from chest X-ray images have the potential to mitigate shortages of radiological expertise and improve clinical efficiency and accuracy (Bannur et al., 2024; Tanida et al., 2023; Chen et al., 2024). However, such models are prone to hallucinations, where the system generates incorrect claims that are not supported by the input image (Liu et al., 2024a; Bai et al., 2024). For example, even high-performing radiology report generation models can hallucinate in approximately 40% of generated sentences (See Table 2 in Appendix B). Such errors can mislead clinicians, with potentially severe consequences for patient care.
@@ 2.1 Uncertainty Modeling and Hallucination Mitigation
Techniques for uncertainty modeling in LLMs fall into three categories based on how much access to the model is required: white-box (using internal components such as intermediate activations), gray-box (using output probabilities), and black-box (using only generated outputs) (Liu et al., 2024b; Huang et al., 2023). Recently, black-box methods measuring self-consistency have emerged as accurate, easy-to-use approaches to assess model output confidence (Zhang et al., 2024; Mündler et al., 2024). Self-Check GPT, a black-box hallucination detection method, compares multiple sampled responses, measures their consistency using tools like BERTscore and LLM prompting, and labels claims that appear inconsistently across samples as hallucinations. Despite using only model outputs, this black-box method demonstrates superior performance compared to gray-box approaches (Manakul et al., 2023). To the best of our knowledge, RadFlag is the first work to extend this concept to VLMs.
In the domain of radiology report generation, various techniques have been explored to mitigate hallucinations and improve accuracy. Existing methods are generally not black-box and often involve fine-tuning a model with new losses or datasets (Banerjee et al., 2024; Anonymous, 2024; Ramesh et al., 2022). Additionally, the intersection of uncertainty modeling and hallucination detection remains relatively unexplored for radiology report generation. To date, only one study has applied uncertainty modeling to remove hallucinations in radiology report generation by using Monte Carlo dropout to estimate uncertainty at both the image and text levels, integrating these estimates into a novel loss function (Wang et al., 2024). While effective, this method relies on access to model weights. In contrast, our sampling-based method provides a versatile, easy-to-use solution that is model-agnostic and compatible with both proprietary and closed-source models.
@@ 3 Methods
The RadFlag method is designed to flag hallucinatory content in generated radiology reports. The method consists of the following key steps:
@@ 3.1 Radiology-Specific Entailment Function
To determine whether a sentence $s$ is entailed by a report $r$, we define a few-shot entailment function tailored for GPT-4 using radiology-specific knowledge. We categorize entailment into three key groups:
@@ 3.2 Entailment Scoring for Candidate Sentences using High-Temperature Reports
We use an entailment score to represent how much support each sentence in the candidate report, which is sampled at a low temperature, has from the corpus of additional reports sampled at a higher temperature. We define this score as $E(s,\{r_{1},\dots,r_{n}\})$ where $s$ is a sentence from the candidate report, and $r_{i}$ is the $i$th report in the corpus of $n$ high-temperature samples. To compute this score for a sentence $s$, we use the entailment function described above to compare $s$ against every high-temperature report, and we then sum the number of reports categorized as “Completely Entailed” or “Partially Entailed.” The final score can range from 0, indicating that no high-temperature report samples support the sentence, to $n$, indicating that all sampled reports support the sentence.
@@ 3.4 Calibrating Sentence-Level Thresholds Using Conformal Risk Control
To label our calibration set, we define a binary entailment function $E_{cal}(s,r_{gt})$. $s$ is a sentence from the candidate report, which was generated by a VLM at a low temperature, and $r_{gt}$ is the corresponding expert-written ground-truth report. As above, we apply our GPT-4 entailment function and treat a sentence as entailed if GPT-4 chooses CE and as non-entailed if GPT-4 chooses PE or NE, indicating any mismatch in content. If a sentence is entailed, $E_{cal}(s,r_{gt})$ = 1; otherwise, $E_{cal}(s,r_{gt})$ = 0.
@@ 3.5 Report-Level Selective Prediction
$f_{rep}(r,\lambda_{2})=\begin{cases}1,&\text{if }h(r)>\lambda_{2}\\ 0,&\text{otherwise}\end{cases}$ (4)
@@ 3.6 Models and Sampling
To evaluate the performance of RadFlag, we apply it to Medversa and RaDialog, two of the highest-performing radiology report generation models on ReXrank, a leaderboard developed for evaluating radiology report generation models from medical images (Lab, 2024; Zhou et al., 2024; Pellegrini et al., 2023). Both Medversa and RaDialog leverage LLMs for report generation, allowing for temperature-based sampling as required by our method. Based on model specifications for ideal hyperparameters, we chose temperature = 0.1 for the low-temperature generation. For the high temperature generations, prior literature referenced a range of 0.9-1.0 as a reasonable temperature setting (Yadkori et al., 2024; Manakul et al., 2023). After qualitatively examining generations from both MedVersa and RaDiolog at various temperatures, we chose temperature 0.5 and 5.0, respectively, since they resulted in diverse generations which were still stylistically coherent.
@@ 4.2 Choosing $\lambda_{2}$ as a Threshold for Flagging Reports
$\lambda_{2}$ is the threshold for flagging reports that have an excessive number of flagged sentences. We explore three thresholds for $\lambda_{2}$ based on the percentage of reports filtered out at each threshold. For both models, $\lambda_{2}$ = 4 provides a conservative threshold, flagging only 5% of all reports in the test set. $\lambda_{2}$ = 3 flags around 10% of all reports in the test set, while $\lambda_{2}$ = 2 flags 25%.
@@ 4.3 Metrics
We assess report quality using four report-level metrics: RadCliQ, RadGraph entity precision, RadGraph entity recall, RadGraph relation precision, and RadGraph relation recall (Yu et al., 2023; Jain et al., 2021). Lower RadCliQ scores indicate better performance, while higher scores indicate better performance on all RadGraph metrics. Of these metrics, RadGraph entity precision serves as the most direct proxy for hallucinations in radiology reports as it measures how many medical entities in the generated report appear in the ground truth. A lower RadGraph entity precision value suggests that hallucinations have introduced unwanted topics into the report.
@@ 4.4 Defining a True Hallucinations Metric
We introduce a custom metric termed “true hallucinations per report” to quantify the occurrence of true hallucinations. We use the same entailment function used to label our calibration set, using GPT-4 to assess whether sentences are supported by their corresponding ground-truth reports or not. To check GPT-4’s assessments, one radiologist and two radiology residents, who are contributors to the paper, labeled 547 sentences as “entailed” or “not entailed” by a ground-truth report. We then compare GPT-4’s results to the clinician labels and find that GPT-4 achieves 84% accuracy. A table with a detailed breakdown is shown in Table 2 of Appendix B.
@@ 5 Results
5.1 RadFlag Flags Hallucinatory Sentences with High Precision
@@
For a more detailed analysis, we categorized each sentence into one of six medical findings: Lungs, Pleura, Cardiomediastinal, Musculoskeletal, Devices, and Other. These categories were developed with input from radiologists (further details in Appendix C).
@@
The effectiveness of hallucination detection varies across different findings categories and can be seen in Figure 3. Both models exhibit strong performance in the “Devices” and “Musculoskeletal” categories, flagging far more hallucinatory than factual statements. However, the sampling method is less effective in the “Pleura,” “Cardiomediastinal,” and “Lungs” categories, where similar numbers of factual and hallucinated sentences are flagged. These differences across categories may reflect underlying variation in how confident and accurate a model is about different topics. This could be because certain features of chest X-ray images are rare in training data and difficult for VLMs to identify, such as device positions or musculoskeletal conditions.
We evaluate the effect of automatically removing all flagged sentences using RadCliQ and RadGraph metrics and found the impact at the report level to be relatively minor. This limited effect is likely due to the small number of flagged sentences, as well as the fact that some factual sentences are incorrectly removed in the process. For details on this analysis, please see Appendix F.
@@ 5.2 RadFlag Flags Lower-Quality, Hallucinatory Reports
Model Quality Threshold $\lambda_{2}$ n Average True Hallucinations per Report RadCliQ v1 RadGraph Entity Precision RadGraph Entity Recall RadGraph Relation Precision RadGraph Relation Recall
@@
We show that RadFlag successfully flags low-quality reports. For Medversa, there is a considerable disparity in the quality of flagged and accepted reports across all metrics and thresholds. Notably, the 57 reports flagged when $\lambda_{2}=2$ have an average of 4.2 hallucinations per report, while the 151 accepted reports have an average of only 1.9 hallucinations per report. We see a similar trend for RaDialog, though the gaps between flagged and accepted reports are less pronounced. According to the true hallucinations metric, RadCliQ, and both RadGraph precision metrics, RaDialog’s accepted reports are higher quality than the flagged reports. In some cases, the flagged reports achieve higher RadGraph recall scores, indicating that these reports contain more hallucinations but also offer better coverage of correct content.
@@ 6 Discussion
While our method shows promise in detecting hallucinations, there is room for further improvement. Our findings suggest that category-specific thresholds ($\lambda$) for flagging errors could improve performance, given observed variations in our method’s performance across categories. Furthermore, complementary methods could be developed to identify other errors such as omissions, providing a more comprehensive assessment of reports. Finally, an implicit assumption in our current method is that AI models exhibit higher confidence when correct and lower confidence when incorrect. Future methods could test how well this assumption holds for specific models, making it possible to identify models that would most benefit from sampling-based methods.
@@ Appendix A Entailment Prompts
The prompts used for our method can be found in Figures 7 and 8. All prompting was done using GPT-4.
@@
Figure 7: Prompt used for entailment score function $E(s,\{r_{1},\dots,r_{n}\})$ to compute entailment score.
@@
Figure 8: Prompt used for ground truth scoring function $g(s,r)$ to compute True Hallucinations. The entailment scheme used in this prompt mirrors the one in Figure 7
@@ Appendix B Clinician Entailment Labels
Table 2 verifies that GPT-4 entailment scores align with clinician labels.
@@
GPT-4
Entailed 279 47
GPT-4
Not Entailed 42 179
Table 2: Clinician entailment labels align well with GPT-4 Scoring Function
@@ Appendix C Category Definitions
We categorize each sentence into one of six medical findings to assess how well RadFlag flags hallucinations in each category. These categories were developed in consultation with experienced radiologists to ensure comprehensive coverage of common chest X-ray observations. Below is a list of keywords that comprise each category, which we use regex to search for when going through each sentence:
@@
  • Pleural Effusion: Pneumothorax, Pneumoth, Effusion, Pleural, Thickening, Scarring, Pleural Mass, Costophrenic, Plaque, Blunting, Blunted, Silhouetting

### v2

@@ Abstract
Generating accurate radiology reports from medical images is a clinically important but challenging task. While current vision language models show promise, they are prone to generating hallucinations, potentially compromising patient care. We introduce RadFlag, a black-box method to enhance the accuracy of radiology report generation. Our method uses a sampling-based flagging technique to find hallucinatory generations that should be removed. We first sample multiple reports at varying temperatures and then use a large language model to identify claims that are not consistently supported across samples, indicating that the model has low confidence in those claims. Using a calibrated threshold, we flag a fraction of these claims as likely hallucinations, which should undergo extra review or be automatically rejected. Our method achieves high precision when identifying both individual hallucinatory sentences and reports that contain hallucinations. As an easy-to-use, black-box system that only requires access to a model’s temperature parameter, RadFlag is compatible with a wide range of radiology report generation models and has the potential to broadly improve the quality of automated radiology reporting.
@@
The primary dataset used in this research is MIMIC-CXR, an open-source chest X-ray report generation dataset publicly available through PhysioNet (Johnson et al., 2019). Code will be made publicly available at https://github.com/rajpurkarlab/RadFlag.
@@ 1 Introduction
Generating accurate radiology reports from medical images is a challenging task that requires a deep understanding of medical imaging and the ability to precisely interpret and communicate complex findings. Medical Vision Language Models (VLMs) that automatically generate radiology reports from chest X-ray images have the potential to mitigate shortages of radiological expertise and improve clinical efficiency and accuracy (Bannur et al., 2024; Tanida et al., 2023; Chen et al., 2024). However, such models are prone to hallucinations, where the system generates incorrect claims that are not supported by the input image (Liu et al., 2024a; Bai et al., 2024). For example, even high-performing radiology report generation models can hallucinate in approximately 40% of generated sentences (see Table 2 in Appendix D). Such errors can mislead clinicians, with potentially severe consequences for patient care.
@@ 2.1 Uncertainty Modeling and Hallucination Mitigation
Techniques for uncertainty modeling in LLMs fall into three categories based on how much access to the model is required: white-box (using internal components such as intermediate activations), gray-box (using output probabilities), and black-box (using only generated outputs) (Liu et al., 2024b; Huang et al., 2023). Recently, black-box methods measuring self-consistency have emerged as accurate, easy-to-use approaches to assess model output confidence (Zhang et al., 2024; Mündler et al., 2024). Self-Check GPT, a black-box hallucination detection method, compares multiple sampled responses, measures their consistency using tools like BERTscore and LLM prompting, and labels claims that appear inconsistently across samples as hallucinations. Despite using only model outputs, this black-box method demonstrates superior performance compared to gray-box approaches (Manakul et al., 2023). Discrete semantic entropy offers another black-box approach that clusters multiple generated outputs based on semantic similarity and computes entropy over these clusters to measure uncertainty (Farquhar et al., 2024). While these methods have proven effective for text-only models, to the best of our knowledge, RadFlag is the first work to extend this concept to VLMs.
In the domain of radiology report generation, various techniques have been explored to mitigate hallucinations and improve accuracy. Existing methods are generally not black-box and often involve fine-tuning a model with new losses or datasets (Banerjee et al., 2024; Anonymous, 2024; Ramesh et al., 2022). Additionally, the intersection of uncertainty modeling and hallucination detection remains relatively unexplored for radiology report generation. To the best of our knowledge, only one study has applied uncertainty modeling to remove hallucinations in radiology report generation by using Monte Carlo dropout to estimate uncertainty at both the image and text levels, integrating these estimates into a novel loss function (Wang et al., 2024). While effective, this method relies on access to model weights. In contrast, our sampling-based method provides a versatile, easy-to-use solution that is model-agnostic and compatible with both proprietary and closed-source models.
@@ 3 Methods
The RadFlag method is designed to flag hallucinatory content in generated radiology reports as depicted in Figure 1. The method consists of the following key steps:
@@ 3.1 Radiology-Specific Entailment Function
To determine whether a sentence $s$ is entailed by a report $r$, we define a few-shot entailment function tailored for GPT-4 using radiology-specific knowledge. The detailed entailment prompt can be found in Appendix A. We categorize entailment into three key groups:
@@ 3.2 Entailment Scoring for Candidate Sentences using High-Temperature Reports
We use an entailment score to represent how much support each sentence in the candidate report, which is sampled at a low temperature, has from the corpus of additional reports sampled at a fixed higher temperature. We define this score as $E(s,\{r_{1},\dots,r_{n}\})$ where $s$ is a sentence from the candidate report, and $r_{i}$ is the $i$th report in the corpus of $n$ high-temperature samples. To compute this score for a sentence $s$, we use the entailment function described above to compare $s$ against every high-temperature report, and we then sum the number of reports categorized as “Completely Entailed” or “Partially Entailed.” The final score can range from 0, indicating that no high-temperature report samples support the sentence, to $n$, indicating that all sampled reports support the sentence.
@@ 3.4 Calibrating Sentence-Level Thresholds Using Conformal Risk Control
To label our calibration set, we define a binary entailment function $E_{cal}(s,r_{gt})$. $s$ is a sentence from the candidate report, which was generated by a VLM at a low temperature, and $r_{gt}$ is the corresponding expert-written ground-truth report. As above, we apply our GPT-4 entailment function and treat a sentence as entailed if GPT-4 chooses CE and as not entailed if GPT-4 chooses PE or NE, indicating any mismatch in content. If a sentence is entailed, $E_{cal}(s,r_{gt})$ = 1; otherwise, $E_{cal}(s,r_{gt})$ = 0.
@@ 3.5 Report-Level Selective Prediction
$f_{rep}(r,\lambda_{2})=\begin{cases}1,&\text{if }h(r)\geq\lambda_{2}\\ 0,&\text{otherwise}\end{cases}$ (4)
@@ 3.6 Models and Sampling
To evaluate the performance of RadFlag, we apply it to Medversa and RaDialog, two of the highest-performing radiology report generation models on ReXrank, a leaderboard developed for evaluating radiology report generation models from medical images (Lab, 2024; Zhou et al., 2024; Pellegrini et al., 2023). Both Medversa and RaDialog leverage LLMs for report generation, allowing for temperature-based sampling as required by our method. Based on model specifications for ideal hyperparameters, we chose temperature = 0.1 for the low-temperature generation. For the high temperature generations, prior literature referenced a range of 0.9-1.0 as a reasonable temperature setting (Yadkori et al., 2024; Manakul et al., 2023). After qualitatively examining generations from both MedVersa and RaDialog at various temperatures, we chose temperature 0.5 and 5.0, respectively, since they resulted in diverse generations which were still stylistically coherent.
@@ 4.2 Choosing $\lambda_{2}$ as a Threshold for Flagging Reports
$\lambda_{2}$ represents the number of flagged sentences in a report as detected by RadFlag. Reports with fewer flagged sentences than $\lambda_{2}$ are kept, while those with $\lambda_{2}$ errors or more are filtered out. This approach results in varying percentages of reports being filtered from the dataset. We explore three thresholds for $\lambda_{2}$. For both models, $\lambda_{2}$ = 4 provides a conservative threshold, flagging only 5% of all reports in the test set. $\lambda_{2}$ = 3 flags around 10% of all reports in the test set, while $\lambda_{2}$ = 2 flags 25%.
@@ 4.3 Metrics
We assess report quality using four report-level metrics: RadCliQ-v1, RadGraph entity precision, RadGraph entity recall, RadGraph relation precision, and RadGraph relation recall (Yu et al., 2023; Jain et al., 2021). Lower RadCliQ-v1 scores indicate better performance, while higher scores indicate better performance on all RadGraph metrics. Of these metrics, RadGraph entity precision serves as the most direct proxy for hallucinations in radiology reports as it measures how many medical entities in the generated report appear in the ground truth. A lower RadGraph entity precision value suggests that hallucinations have introduced unwanted topics into the report.
@@ 4.4 Defining a True Hallucinations Metric
We introduce a custom metric termed “true hallucinations per report” to quantify the occurrence of true hallucinations. We use the same entailment function used to label our calibration set, using GPT-4 to assess whether sentences are supported by their corresponding ground-truth reports or not. To check GPT-4’s assessments, one radiologist and one senior internal medicine resident, who are listed in the acknowledgements, and one clinician, who is a contributor to the paper, labeled 547 sentences as “entailed” or “not entailed” by a ground-truth report. We then compare GPT-4’s results to the clinician labels and find that GPT-4 achieves 84% accuracy. A table with a detailed breakdown is shown in Appendix B.
@@ 5 Results
5.1 RadFlag Flags Hallucinatory Sentences with High Precision
@@
For a more detailed analysis, we categorized each sentence into one of six medical findings: Lungs, Pleural, Cardiomediastinal, Musculoskeletal, Devices, and Other. These categories were developed with input from radiologists (further details in Appendix C).
@@
The effectiveness of hallucination detection varies across different findings categories and can be seen in Figure 3. Detailed category counts are provided in Appendix E. Both models exhibit strong performance in the “Devices” and “Musculoskeletal” categories, flagging far more hallucinatory than factual statements. However, the sampling method is less effective in the “Pleural,” “Cardiomediastinal,” and “Lungs” categories, where similar numbers of factual and hallucinated sentences are flagged. These differences across categories may reflect underlying variation in how confident and accurate a model is about different topics. This could be because certain features of chest X-ray images are rare in training data and difficult for VLMs to identify, such as device positions or musculoskeletal conditions.
We evaluate the effect of automatically removing all flagged sentences using RadCliQ-v1 and RadGraph metrics and found the impact on report-level metrics to be relatively minor. This limited effect is likely due to the small number of flagged sentences, as well as the fact that some factual sentences are incorrectly removed in the process. For details on this analysis, please see Appendix F.
@@ 5.2 RadFlag Flags Lower-Quality, Hallucinatory Reports
Model Quality Threshold $\lambda_{2}$ n Average True Hallucinations per Report RadCliQ-v1 RadGraph Entity Precision RadGraph Entity Recall RadGraph Relation Precision RadGraph Relation Recall
@@
Table 1 shows that RadFlag successfully flags low-quality reports. For Medversa, there is a considerable disparity in the quality of flagged and accepted reports across all metrics and thresholds. Notably, the 57 reports flagged when $\lambda_{2}=2$ have an average of 4.2 hallucinations per report, while the 151 accepted reports have an average of only 1.9 hallucinations per report. We see a similar trend for RaDialog, though the gaps between flagged and accepted reports are less pronounced. Figure 5 shows the distribution of hallucinations in flagged versus accepted reports across both models. According to the true hallucinations metric, RadCliQ-v1, and both RadGraph precision metrics, RaDialog’s accepted reports are higher-quality than the flagged reports. In some cases, the flagged reports achieve higher RadGraph recall scores, indicating that these reports contain more hallucinations but also offer better coverage of correct content.
Our approach also achieves more reliable differentiation between high and low-quality reports, outperforming baseline entropy methods (see Appendix G). Despite requiring access to per-token probabilities and therefore requiring more technical expertise to implement, we found entropy baselines were not successful in identifying hallucinatory outputs. This is likely because there are many valid ways to express the same radiological idea, thus low probability does not equate to low confidence in the overall claim.
@@ 6 Discussion
While our method shows promise in detecting hallucinations, there is room for further improvement. We acknowledge that our reliance on GPT-4 for both entailment scoring and hallucination detection may introduce correlated errors that could inflate our accuracy metrics. While we validated GPT-4’s entailment scores with clinicians, future work would benefit from more extensive clinical validation of the hallucination detection process itself. Our findings suggest that category-specific thresholds ($\lambda_{1}$) for flagging errors could improve performance, given observed variations in our method’s performance across categories. Furthermore, complementary methods could be developed to identify other errors such as omissions, providing a more comprehensive assessment of reports. Finally, an implicit assumption in our current method is that AI models exhibit higher confidence when correct and lower confidence when incorrect. Future methods could test how well this assumption holds for specific models, making it possible to identify models that would most benefit from sampling-based methods.
Acknowledgments
We would like to thank Dr. Sara Gomez Villegas and Dr. Jung Oh Lee for their clinical expertise and help in evaluating GPT-4’s performance through entailment labeling.
@@ Appendix A Entailment Prompts
The prompts used for our method can be found in Figures A1 and A2. All prompting was done using GPT-4.
@@
Figure A1: Prompt used for entailment score function $E(s,\{r_{1},\dots,r_{n}\})$ to compute entailment score.
@@
Figure A2: Prompt used for ground truth scoring function $g(s,r)$ to compute True Hallucinations. The entailment scheme used in this prompt mirrors the one in Figure A1
@@ Appendix B Clinician Entailment Labels
Table A1 verifies that GPT-4 entailment scores align with clinician labels.
@@
GPT-4 Entailed 279 47
GPT-4 Not Entailed 42 179
Table A1: Clinician entailment labels align well with our GPT-4 Scoring Function.
@@ Appendix C Category Definitions
We categorize each sentence into one of six medical findings to assess how well RadFlag flags hallucinations in each category. These categories were developed in consultation with experienced radiologists to ensure comprehensive coverage of common chest X-ray observations. Below is a list of keywords that comprise each category. If sentences contain keywords from multiple categories, we assign them to the first matching category based on the following priority order: Devices, Cardiomediastinal, Lungs, Musculoskeletal, Pleura, and Other. We use regular expressions to search for these keywords when going through each sentence:
@@
  • Pleural: Pneumothorax, Pneumoth, Effusion, Pleural, Thickening, Scarring, Pleural Mass, Costophrenic, Plaque, Blunting, Blunted, Silhouetting